OpenStudy (anonymous):

e^x=-x+1, how do you solve this? @MIT 18.01 Single …

OpenStudy (jamesj):

In this case x = 0 is the solution.

OpenStudy (jamesj):

Normally with an expression e^(ax) = polynomial in x, numerical methods are needed. But with yours we can see a solution by inspection.

OpenStudy (anonymous):

friend it is easy:just draw the graphs of y=e^x & y=-x+1....>the intersection point is the solution....this is a transcendental equation